# Reliable Simulation of Quantum Channels: the Error Exponent

### v4

@@ III Symmetrization and de Finetti Reduction
Proof. For any pure state $\varphi_{RA^{n}}$, we have
@@ V-A Channel Simulation with a Fixed Input
Proof. Let $U_{\mathcal{N}}^{A\rightarrow BE}$ be the Stinespring dilation of the channel $\mathcal{N}_{A\rightarrow B}$. We start with an introduction of quantum state merging with respect to the state $\varphi_{REB}={\operatorname{id}}_{R}\otimes U_{\mathcal{N}}^{A\rightarrow BE}(\varphi_{RA})$. Suppose that $\varphi_{REB}$ is held by a referee ($R$), Alice ($E$), and Bob ($B$). This is compactly represented by $\varphi_{R:E:B}$. That is, when multiple systems are separated by two colons, it means that the three parts from left to right are respectively held by the referee, Alice and Bob. Based on the protocol introduced in [51, 50], a tight achievability bound has been derived in [33], for merging the $B$ part of $\varphi_{R:E:B}$ to Alice. The merging scheme consists of the following steps:
@@
Proof. Because $\zeta_{A^{n}}$ is invariant under permutations, for any $\pi\in S_{n}$ the pure states $W^{\pi}_{A^{n}}(\zeta_{R_{n}A^{n}})$ and $\zeta_{R_{n}A^{n}}$ are identical on the $A^{n}$ system. So, there is a unitary operation $V_{R_{n}}$ such that
@@
Proof of Theorem 9. Since we can always symmetrize the simulation channel, according to Lemma 8, it suffices to consider the de Finetti input state $\zeta_{R_{n}A^{n}}$. Let $\zeta_{R_{n}B^{n}}=\mathcal{N}_{A\rightarrow B}^{\otimes n}(\zeta_{R_{n}A^{n}})$. From now on, we fix an $s\in(0,1]$. As shown in Lemma 14, we can choose a symmetric (permutation-invariant) state $\sigma^{*}_{B^{n}}$ to minimize $D_{1+s}(\zeta_{R_{n}B^{n}}\|\zeta_{R_{n}}\otimes\sigma_{B^{n}})$, such that
@@
Proof of Theorem 11: Eq. (3). It is a straightforward consequence of Theorem 9 that
@@ VI-A One-Shot Converse Bound
Proof. It is straightforward to check that
@@
Proof. Let $\mathcal{M}_{A\rightarrow B}$ be a general reverse Shannon simulation with not more than $c$ bits of classical communication. It consists of (a) using a shared bipartite entangled state $\omega_{A_{1}B_{1}}$ for assistance, (b) applying local operation $\mathcal{E}_{AA_{1}\rightarrow X}$ with $\log|X|\leq c$ at Alice’s side, (c) sending $X$ from Alice to Bob via a noiseless classical channel $\overline{{\operatorname{id}}}_{X\rightarrow X}:\ \rho_{X}\mapsto\sum_{x}\langle x|\rho_{X}|x\rangle|x\rangle\!\langle x|_{X}$ with $\{|x\rangle\}_{x}$ a standard basis of $\mathcal{H}_{X}$, (d) applying local operation $\mathcal{D}_{XB_{1}\rightarrow B}$ at Bob’s side.
@@
Proof of Theorem 10. The proof is done by combining Lemma 15 and Lemma 16. ∎
@@ VI-B Final Proof of the Reliability Function
Proof. At first, we prove the concavity. Let $\rho_{A},\sigma_{A}\in\mathcal{S}(A)$ be density matrices, and let $\psi_{RA}$ and $\phi_{RA}$ be the purifications of $\rho_{A}$ and $\sigma_{A}$, respectively. For any $0\leq\lambda\leq 1$, consider the convex combination $\lambda\rho_{A}+(1-\lambda)\sigma_{A}$. By introducing another system $F$, its purification can be written as
@@
Proof. It is shown in [48] that $s\mapsto I_{1+s}(\mathcal{N}_{A\rightarrow B})$ is monotonically increasing. This implies that $s\mapsto sI_{1+s}(\mathcal{N}_{A\rightarrow B})$ is monotonically increasing too. To see that $s\mapsto sI_{1+s}(\mathcal{N}_{A\rightarrow B})$ is convex, we write $sI_{1+s}(\mathcal{N}_{A\rightarrow B})=\max\limits_{\rho_{A}}\{sI_{1+s}(\mathcal{N}_{A\rightarrow B},\rho_{A})\}$ and then note that $s\mapsto sI_{1+s}(\mathcal{N}_{A\rightarrow B},\rho_{A})$ is convex by Ref. [55]. ∎
Proof of Theorem 11: Eqs. (4) and (5). To prove Eq. (4), we employ Theorem 10. In Ref. [33], it is shown that for any state $\rho_{RB}$ and $r\geq 0$,
@@ VII Conclusion and Discussion
When the classical communication rate is above the critical point, we are unable to obtain the reliability function. In fact, the existence of a critical point, of which at one side the problem remains unsolved, is a common phenomenon in the study of reliability functions [19, 24, 32]. Works in literature (e.g., [19, 58, 59]) suggest that combinatorics may come into play. We leave it as an open problem. We note that in the independent work [39], the moderate deviation expansion for reverse Shannon simulation of quantum channels in the low-error case has been derived. The second-order asymptotics of this task remains a major open problem. Yet another interesting problem related to our work, is to find the reliability function for entanglement-assisted communication over quantum channels.
@@
Proof of Lemma 8. We start with an arbitrary state $\varphi_{TA^{n}}$ in which $T$ is any reference system. Set
@@
Proof. (1) Since $s\mapsto sI_{1+s}(\mathcal{N})$ is convex on $[0,\infty)$ by Lemma 18, the function $f(s)=s(r-I_{1+s}(\mathcal{N}))$ is concave. When $f^{\prime}(1)>0$, which is equivalent to $r>R_{\rm{critical}}$, we have $E_{l}(r)=\frac{1}{2}f(1)=\frac{1}{2}(r-I_{2}(\mathcal{N}))$. (2) The function $s\mapsto I_{1+s}(\mathcal{N})$ is monotonically increasing [48] and obviously bounded for $s\in[0,\infty)$. So, the limit $\lim\limits_{s\rightarrow\infty}I_{1+s}(\mathcal{N})$ exists and the statement follows. ∎

### v5

@@ III Symmetrization and de Finetti Reduction
Proof:
For any pure state $\varphi_{RA^{n}}$, we have
@@ IV Main Results
To see at what range of the communication rate we have obtained the exact error exponent, we computer the critical value and the mutual information of depolarizing channels; see Example 20 in the Appendix. It shows that $R_{\text{critical}}$ is considerably larger than $I(\mathcal{N})$.
@@ V-A Channel Simulation with a Fixed Input
Proof:
Let $U_{\mathcal{N}}^{A\rightarrow BE}$ be the Stinespring dilation of the channel $\mathcal{N}_{A\rightarrow B}$. We start with an introduction of quantum state merging with respect to the state $\varphi_{REB}={\operatorname{id}}_{R}\otimes U_{\mathcal{N}}^{A\rightarrow BE}(\varphi_{RA})$. Suppose that $\varphi_{REB}$ is held by a referee ($R$), Alice ($E$), and Bob ($B$). This is compactly represented by $\varphi_{R:E:B}$. That is, when multiple systems are separated by two colons, it means that the three parts from left to right are respectively held by the referee, Alice and Bob. Based on the protocol introduced in [51, 50], a tight achievability bound has been derived in [33], for merging the $B$ part of $\varphi_{R:E:B}$ to Alice. The merging scheme consists of the following steps:
@@
Proof:
Because $\zeta_{A^{n}}$ is invariant under permutations, for any $\pi\in S_{n}$ the pure states $W^{\pi}_{A^{n}}(\zeta_{R_{n}A^{n}})$ and $\zeta_{R_{n}A^{n}}$ are identical on the $A^{n}$ system. So, there is a unitary operation $V_{R_{n}}$ such that
@@
Proof:
Since we can always symmetrize the simulation channel, according to Lemma 8, it suffices to consider the de Finetti input state $\zeta_{R_{n}A^{n}}$. Let $\zeta_{R_{n}B^{n}}=\mathcal{N}_{A\rightarrow B}^{\otimes n}(\zeta_{R_{n}A^{n}})$. From now on, we fix an $s\in(0,1]$. As shown in Lemma 14, we can choose a symmetric (permutation-invariant) state $\sigma^{*}_{B^{n}}$ to minimize $D_{1+s}(\zeta_{R_{n}B^{n}}\|\zeta_{R_{n}}\otimes\sigma_{B^{n}})$, such that
@@
Proof:
It is a straightforward consequence of Theorem 9 that
@@ VI-A One-Shot Converse Bound
Proof:
It is straightforward to check that
@@
Proof:
Let $\mathcal{M}_{A\rightarrow B}$ be a general reverse Shannon simulation with not more than $c$ bits of classical communication. It consists of (a) using a shared bipartite entangled state $\omega_{A_{1}B_{1}}$ for assistance, (b) applying local operation $\mathcal{E}_{AA_{1}\rightarrow X}$ with $\log|X|\leq c$ at Alice’s side, (c) sending $X$ from Alice to Bob via a noiseless classical channel $\overline{{\operatorname{id}}}_{X\rightarrow X}:\ \rho_{X}\mapsto\sum_{x}\langle x|\rho_{X}|x\rangle|x\rangle\!\langle x|_{X}$ with $\{|x\rangle\}_{x}$ a standard basis of $\mathcal{H}_{X}$, (d) applying local operation $\mathcal{D}_{XB_{1}\rightarrow B}$ at Bob’s side.
@@
Proof:
The proof is done by combining Lemma 15 and Lemma 16. ∎
@@ VI-B Final Proof of the Reliability Function
Proof:
At first, we prove the concavity. Let $\rho_{A},\sigma_{A}\in\mathcal{S}(A)$ be density matrices, and let $\psi_{RA}$ and $\phi_{RA}$ be the purifications of $\rho_{A}$ and $\sigma_{A}$, respectively. For any $0\leq\lambda\leq 1$, consider the convex combination $\lambda\rho_{A}+(1-\lambda)\sigma_{A}$. By introducing another system $F$, its purification can be written as
@@
Proof:
It is shown in [48] that $s\mapsto I_{1+s}(\mathcal{N}_{A\rightarrow B})$ is monotonically increasing. This implies that $s\mapsto sI_{1+s}(\mathcal{N}_{A\rightarrow B})$ is monotonically increasing too. To see that $s\mapsto sI_{1+s}(\mathcal{N}_{A\rightarrow B})$ is convex, we write $sI_{1+s}(\mathcal{N}_{A\rightarrow B})=\max\limits_{\rho_{A}}\{sI_{1+s}(\mathcal{N}_{A\rightarrow B},\rho_{A})\}$ and then note that $s\mapsto sI_{1+s}(\mathcal{N}_{A\rightarrow B},\rho_{A})$ is convex by Ref. [55]. ∎
Proof:
To prove Eq. (4), we employ Theorem 10. In Ref. [33], it is shown that for any state $\rho_{RB}$ and $r\geq 0$,
@@ VII Conclusion and Discussion
When the classical communication rate is above the critical point, we are unable to obtain the reliability function. In fact, the existence of a critical point, of which at one side the problem remains unsolved, is a common phenomenon in the study of reliability functions [19, 24, 32]. Works in literature (e.g., [19, 58, 59]) shows that combinatorics come into play for channel communication. We do not know whether this is still the case for our simulation setting. We leave it as an open problem. We note that in the independent work [39], the moderate deviation expansion for reverse Shannon simulation of quantum channels in the low-error case has been derived. The second-order asymptotics of this task remains a major open problem. Yet another interesting problem related to our work, is to find the reliability function for entanglement-assisted communication over quantum channels.
Acknowledgements
The authors are grateful to one of the anonymous referees for suggesting look at the gap between the critical point and the channel’s mutual information, which leads to Example 20. They also would like to thank Yimeng Cao for his assistance in drawing Fig. 2.
@@ Acknowledgements
Proof:
We start with an arbitrary state $\varphi_{TA^{n}}$ in which $T$ is any reference system. Set
@@
Proof:
(1) Since $s\mapsto sI_{1+s}(\mathcal{N})$ is convex on $[0,\infty)$ by Lemma 18, the function $f(s)=s(r-I_{1+s}(\mathcal{N}))$ is concave. When $f^{\prime}(1)>0$, which is equivalent to $r>R_{\rm{critical}}$, we have $E_{l}(r)=\frac{1}{2}f(1)=\frac{1}{2}(r-I_{2}(\mathcal{N}))$. (2) The function $s\mapsto I_{1+s}(\mathcal{N})$ is monotonically increasing [48] and obviously bounded for $s\in[0,\infty)$. So, the limit $\lim\limits_{s\rightarrow\infty}I_{1+s}(\mathcal{N})$ exists and the statement follows. ∎
Figure 2: Critical value and mutual information of the qubit depolarizing channel. Upper curve is $R_{\rm{critical}}(\mathcal{N}^{(p)})$, and lower curve is $I(\mathcal{N}^{(p)})$.
In the following example, we calculate the mutually information and the critical value for the depolarizing channels.
Example 20.
We consider the depolarizing channels $\mathcal{N}_{A\rightarrow B}^{(p)}$ with $|A|=|B|=d$ and $p\in[0,1]$, given by
$\mathcal{N}^{(p)}(\rho):=(1-p)\rho+p\frac{\mathbbm{1}}{d}.$
Due to the symmetry of this channel, namely, its being covariant with respect to the Heisenberg-Weyl group, we can easily verify that for $s\geq 0$,
$I_{1+s}(\mathcal{N}^{(p)})=D_{1+s}\left(p\Phi_{AB}+(1-p)\frac{\mathbbm{1}_{A}}{d}\otimes\frac{\mathbbm{1}_{B}}{d}\Big{\|}\frac{\mathbbm{1}_{A}}{d}\otimes\frac{\mathbbm{1}_{B}}{d}\right),$
where $\Phi_{AB}$ is a maximally entangled pure state. With this, we get by direct computation that
$I(\mathcal{N}^{(p)})$ $=\frac{d^{2}p+1-p}{d^{2}}\log(d^{2}p+1-p)+\frac{(d^{2}-1)(1-p)}{d^{2}}\log(1-p),$
$R_{\rm{critical}}(\mathcal{N}^{(p)})$ $=\frac{(d^{2}p+1-p)^{2}\log(d^{2}p+1-p)}{(d^{2}p+1-p)^{2}+(d^{2}-1)(1-p)^{2}}+\frac{(d^{2}-1)(1-p)^{2}\log(1-p)}{(d^{2}p+1-p)^{2}+(d^{2}-1)(1-p)^{2}}.$
In Fig. 2, we depict these two quantities as functions of $p$, in the special qubit case $d=2$. It shows that $R_{\rm{critical}}(\mathcal{N}^{(p)})$ is considerably larger than $I(\mathcal{N}^{(p)})$. On the other hand, when $d$ is big, $R_{\rm{critical}}(\mathcal{N}^{(p)})$ can be much larger than $I(\mathcal{N}^{(p)})$, since for constant $p$ we have
$I(\mathcal{N}^{(p)})$ $\sim 2p\log d,$
$R_{\rm{critical}}(\mathcal{N}^{(p)})$ $\sim 2\log d.$